# ConfusionPrompt: Practical Private Inference for Online Large Language Models

### v4

@@ 5.3.5 Local Model Selection
In this section we justify the selection of decomposor and recomposor. For decomposor, we obtain the BLEU between the generated decomposition and reference provided by the dataset. For recomposor, the models are evaluated in terms of the accuracy metrics described in Section 5.3.2, using the golden decomposition and sub-answer as training and testing inputs.
Table 4 presents the experiment results on MuSiQue using models from three different classes of varying sizes. While Flan-T5-large exhibits the best performance for both tasks, we select BART-large as the decomposition and recomposition model due to its balanced trade-off between model size and performance.
@@
Parameter size Decomposer Recomposer
@@
BART-base 139M 0.314 0.563 0.564 0.446
BART-large 406M 0.402 0.640 0.641 0.525
T5-base 223M 0.343 0.581 0.580 0.484
T5-large 737M 0.408 0.646 0.645 0.554
Flan-T5-base 248M 0.367 0.608 0.609 0.518
Flan-T5-large 783M 0.458 0.671 0.671 0.577
@@ 6 Discussion
In this section, we provide further clarifications and discussions on practical issues of our framework.
@@
Protection capability of privacy model. Our $(\lambda,\mu,\rho)$-privacy model establishes requirements for fabricated prompts to mitigate privacy risks. Rather than offering a worst-case privacy guarantee such as LDP, the model quantifies the criteria that a group of prompts should satisfy. Our empirical experiments associate the privacy parameters with attack success rate. It is important to note that the privacy model should be regarded as an enhancement over the insecure status quo, rather than a substitute for formal privacy guarantees.
@@
Tradeoff between privacy and query cost. Whereas traditional perturbation-based methods trade privacy for utility, our ConfusionPrompt trades privacy for query complexity while providing accurate responses. This approach is particularly suitable for clients with high privacy demands who can accommodate these costs, such as finance and healthcare organizations. Additionally, although query costs rise linearly with the number of tokens in API services, users of chatbot services can submit unlimited requests under a fixed subscription fee, allowing for enhanced privacy protection at a manageable cost.
@@ 7 Conclusion
This paper proposes a private inference framework on online LLMs, termed ConfusionPrompt. We deploy three local models on the user side: (i) decomposer that maps an original prompt to a sequence of sub-prompts, (ii) generator that produces pseudo prompts by replacing private attributes in the genuine sub-prompts, (iii) recomposer that maps the decomposed prompt-response pairs from the cloud service to the final response. Such design endows our framework with advantages over previous protocols that: (i) it can be seamlessly integrated with existing black-box LLMs where the service providers have no need to modify their already-built framework, and (ii) it achieves better privacy-utility trade-off than existing LDP-based privatization methods. We develop a $(\lambda,\mu,\rho)$-privacy model to formulate the requirement for a privacy-preserving group of prompts, and accordingly provide a complexity analysis to demonstrate the benefits of the decomposition module. Experiments show that our ConfusionPrompt significantly outperforms the inference with local open-sourced LLMs in terms of utility and computation cost.
Appendix A Proof of Theorem 11 and 12
We begin with the proof for Theorem 11 as followed:
Proof.
For the prompt with single paragraph, the significance can by represented as:
$\begin{gathered}\mathrm{Sig}(U(p_{0}),\boldsymbol{P})=\max_{u_{k}}\max_{\mathcal{V}}\frac{\sum_{i=1}^{n}\bar{H}(\mathcal{V},p_{0},p_{i})}{\sum_{i=1}^{n}\bar{H}(\mathcal{V}\backslash u_{k},p_{0},p_{i})}\\ {\rm s.t.}u_{k}\in\mathcal{V}\subseteq U(p_{0}))\end{gathered}$ (11)
, where $p_{0}$ is the genuine prompt, and $p_{i}$ is the $i^{\mathrm{th}}$ prompt in the prompt group, and:
$\bar{H}(\mathcal{V},p_{0},p_{i})=\left\{\begin{array}[]{ll}1&\mathrm{Corr}\left(u,p_{0},p_{i}\right)=u\ \forall u\in\mathcal{V}\\ &\mathrm{or}\ \mathcal{V}=\emptyset\\ 0&\rm{otherwise}\end{array}.\right.$ (12)
For $\mathrm{Sig}(U(p_{0}),\boldsymbol{P})=\mu$, it’s obvious that there should be at least $1/\mu$ distinct values, including genuine and fake values, assigned to each attributes. Otherwise suppose $u_{k}$ have less than $1/\mu$ distinct values, we would have:
$\frac{\sum_{j=1}^{|\boldsymbol{P}|}H(u_{k},p_{0},p_{j})}{|\boldsymbol{P}|}>\mu,$ (13)
which violates the expression 11.
Next, we claim that each possible combination of attribute set values must appear at least once within the prompt group. To proof by contradiction, let’s suppose that there is one combination missing in the group while the others appear once. Then we can consider two cases:
(a) The missing combination doesn’t contain any true values. Then the significance of other single attribute must be larger than $\mu$. This is because for a set containing all combinations, the proportion of the occurrence for each attribute is exactly $\mu$. Therefore, such removal would make their significance smaller.
(b) The missing combination contains a subset of true values $\mathcal{V}\subseteq U(p_{0})$. Then conditioning on the subset $\mathcal{V}$, the proportion of any other single attribute would be over $\mu$, as their original proportion is $\mu$. Therefore, for any $u_{k}\notin\mathcal{V}$, it holds that:
$\frac{\sum_{i=1}^{n}\bar{H}(\mathcal{V}\cup u_{k},p_{0},p_{i})}{\sum_{i=1}^{n}\bar{H}(\mathcal{V},p_{0},p_{i})}>\mu,$ (14)
which violates the expression 11.
To make each combination occur at least once, we should have at least $(1/\mu)^{|U(p_{0})|}$ prompts in the group. ∎
Following that, we can proceed to the proof for Theorem 12:
Proof.
Let $\mu=\mathrm{Sig}(U(p_{0}),\boldsymbol{P})$ denotes the significance. According to Theorem 11, for each sub-query $p_{0h}\in\boldsymbol{p}_{0}$, we should have at least $(1/\mu)^{\|U(p_{0h})\|}$ prompts to satisfy:
$\begin{gathered}\max_{u_{k}}\max_{\mathcal{V}}\frac{\sum_{i=1}^{n}\bar{H}(\mathcal{V},p_{0h},p_{ih})}{\sum_{i=1}^{n}\bar{H}(\mathcal{V}\backslash u_{k},p_{0h},p_{ih})}\leq\mu\\ {\rm s.t.}u_{k}\in\mathcal{V}\subseteq U(p_{0h})\end{gathered}$ (15)
Then the proof completes by summing up the number of prompts over all sub-queries. ∎
Appendix B Bound on Inference Attack
In this section, we consider the scenario where an attacker attempts to identify the target attribute by prior knowledge of other attributes and the combination pattern of attribute values. The following theorem states that an curious server can do no better than random guessing if the combination of values are uniformly distributed:
Lemma 1.
Suppose the attacker is guessing the attributes from attribute combination pattern in $\boldsymbol{P}$. The attacker can do no more better than random guessing if the following condition is satisfied:
  • 1.
    Each combination of attribute set has the same occurrence times in the prompt groups $\boldsymbol{P}$, i.e., the combinations are uniformly distributed.
From the proof of Theorem 11 and 12, we can construct a uniformly distributed prompt group under the optimal number of prompts. Therefore, we can turn to the success rate under an random guessing attack, which is bounded by Lemma 2.
Lemma 2.
Given a group of prompts $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ with $(\lambda,\mu,\rho)$-privacy, suppose that the attacker have prior knowledge about the attribute subset $\mathcal{V}\subseteq U(\boldsymbol{p}_{0})$. By random guessing, the probability of correctly identifying any target attribute $u\in U(\boldsymbol{p}_{0})\land u\notin\mathcal{V}$ is upper bounded by $\mu$.
Proof.
According to Definition 8, it holds that:
$\begin{gathered}\frac{\sum_{j=1}^{n}\bar{H}(\mathcal{V},p_{0h},p_{jh})}{\sum_{j=1}^{n}\bar{H}(\mathcal{V}\backslash u,p_{0h},p_{jh})}\leq\mu\end{gathered}$ (16)
for any $h$ and $u$. In other words, the proportion of any attribute $u$ conditioning on the prior knowledge attribute set $\mathcal{V}$ is upper bounded by $\mu$. Therefore, by random guessing, the probability of correctly identifying any target attribute $u$ is upper bounded by $\mu$. ∎
An interpretation of the above lemmas is that under $(\lambda,\mu,\rho)$-privacy, we can always construct a group of prompts with optimal complexity, where the probability of correctly identifying any target attribute is upper bounded by $\mu$, if the attacker utilizes the combination pattern of attributes (ignoring the semantic meaning).
Appendix C Experiment
C.1 Construction of Private Attributes
During step 1, we label the private attributes according to the following guidelines:
  • 1.
    Identify the key components for a query, including but not limited to proper nouns, phrases, verbs, and adjectives.
  • 2.
    Ensure that there is no overlapping information between different private attributes within the same query.
  • 3.
    Each private attribute should be as succinct as possible. For example, instead of labeling ”spouse of the Green performer” as a single attribute, it is preferable to split it into ”spouse” and ”Green performer” as separate attributes.
The third point is to ensure privacy protection even when an attacker possesses prior knowledge of certain attributes. Table 5 provides several examples of private attributes.
Table 5: Examples of private attributes in StrategyQA and MuSiQue.
Query Private Attributes
Are blue lips normal? blue lips, normal
Could ten gallons of seawater crush a six year old? seawater crush, ten gallons, six year old
Who is the spouse of the Green performer? spouse, Green performer
What instrument is played by the person from The Blackout All-Stars? play instrument, The Blackout All-Stars
What is the capital of the county that Pine Springs is located in? capital of the county, Pine Springs
Table 6: Sample Questions and Associated Risks for Each Category of PrivCP.
Type N Example Question Potential Risks
Business 20 ‘As a VP of marketing at MovieMagic, I am considering a shift in our advertising spend from TV to digital platforms. What are the key factors to consider?’ Reveal the strategic considerations of the company
Legal 20 ‘I have a meeting with a client at Clark Legal Services in Boston who wants to create a trust. What are the different types of trusts available?’ Reveal the sensitive case details of a client
Health 20 ‘My 3-year-old daughter, Emily, has just been diagnosed with a peanut allergy. How can I ensure she avoids peanuts at her daycare?’ Reveal the medical and personal info of family members
Career 20 ‘I have worked as a civil engineer at AECOM for five years in Los Angeles. What certifications should I pursue for career advancement?’ Reveal the personal career experience and work location
Education 20 ‘I am a parent evaluating private schools in Chicago for my 6-year-old son, Matthew. What are the key factors to consider in choosing a school with special needs support?’ Reveal the personal info and
Social 20 ‘My company is hosting a corporate event at the Hilton in Las Vegas. How can I network effectively with potential clients?’ Reveal the location and purpose of the private business event
Personal 20 ‘I am moving to a penthouse in Manhattan in July. How can I best use natural light in my home decor?’ Reveal the residential location, moving schedule and personal preference for decor
In step 2, we evaluate the generative quality of various LLMs on five models: i) GPT-4-Turbo with few-shot examples, ii) GPT-3.5-Turbo finetuned with 100 samples, iii) BART-large [44], iv) T5-large [45], and v) Flan-T5-Xlarge [46]. The latter three open-source models are finetuned with 1000 samples.
Table 7 presents the performance on a human-labeled test dataset. It can be observed that the finetuned version of GPT-3.5-turbo gives the best result, and thus it’s used as the attribute extraction model throughout our experiment.
Table 7: Comparison of various attribute extraction models.
Method StrategyQA MuSiQue
F1 roughL EM F1 roughL EM
GPT-4-Turbo 0.689 0.748 0.080 0.811 0.762 0.234
GPT-3.5-Turbo 0.803 0.881 0.421 0.846 0.789 0.333
BART-large 0.713 0.690 0.227 0.664 0.634 0.118
T5-large 0.603 0.595 0.102 0.381 0.383 0.010
Flan-T5-large 0.669 0.683 0.243 0.380 0.383 0.026
C.2 Training of local models
In this section, we describe the training of decomposer, generator, and recomposer, respectively.
Training of decomposer: The decomposition model is finetuned on a pretrained BART model [44] with 406M parameters. In the first stage, we utilize the decomposition samples from StrategyQA and MuSiQue as demonstration data. In the second stage, we evaluate the response through the following steps: (i) extract the sub-answers using gpt-4-turbo and obtain the final answers using a recomposition model to judge whether the decomposition returns a correct answer; (ii) compute the complexity measured by the number of required pseudo prompts under $\mu=10$.
During step (i), to address the recomposer’s limitations in determining the correctness of the final answer, we train a robust recomposition model by fine-tuning LLaMa2-7B. Note that during inference, we adopt a much smaller recomposition model for efficiency.
Training of generator: The generator is finetuned on the BART-large model. To collect demonstration data for the first stage, we prompt GPT-4-turbo to generate multiple pseudo prompts given the raw prompt and its private attributes using the following template. {mdframed}[linecolor=red!50!black, backgroundcolor=yellow!20]    Please replace the phrases {attributes} in the each of the following sentences, such that each sentence is fluent and reasonable, and the alterntive phrases have irrelevant meaning as {attributes}.
Please return {# of replaces} replacements for each sentence.
Strictly respond in the form of JSON with the following format: {”S1”: [”replacement 1”, ”replacement 2”,…], ”S2”: [”replacement 1”, ”replacement 2”,…]}.
Sentences: {dictionary of sentences}
In the second, we employ two models to evaluate the semantic similarity and genuineness, including a local similarity evaluation model and discriminator. For similarity evaluation model, we adopt a fine-tuned version of MiniLM-6L model [47] to extract the embedding of each private attribute. The semantic relevance between a pair of attributes is given by the cosine similarity between their embeddings. For discriminator, We leverage GPT-4o to construct the training dataset for training a local discriminator. As the genuineness is closely related to the sentence’s fluency, we instruct GPT-4o to evaluate the fluency for each sentence [48] (see detailed instruction in C.3).
Training of recomposer: The recomposition model is finetuned on a pretrained RoBERTa [49] and BART-large [44] model, respectively, for strategyQA and MuSiQue. We utilize the sub-prompts & sub-responses given in StrategyQA and MuSiQue dataset as well as the final response to train a recomposition model. To improve the performance of recomposor, we pretrain the model on two addition datasets, SQuAD [50] and DROP [51].
C.3 Semantic Similarity Model and Discriminator
The comparison data collection for the generator involves a local similarity evaluation model and discriminator.
Similarity evaluation model: We adopt a finetuned version of MiniLM-6L model [47] to extract the embedding of each private attribute. The semantic relevance between a pair of attributes is given by the cosine similarity between their embeddings.
Discriminator: We leverage GPT-4 to construct the training dataset for training a local discriminator. As the genuineness is closely related to the sentence’s fluency, we prompt GPT-4 to evaluate the fluency for each sentence with the following template [48]: {mdframed}[linecolor=red!50!black, backgroundcolor=yellow!20]    Given multiple sentences, use the scoring rules below to score each sentence’s fluency on a scale of 1 to 4:
1. Score 1: Incomprehensible. Inarticulate/ non-fluent sentence.
2. Score 2: Low Quality. Partially fluent sentence: (a) only half of the sentence is fluent or (b) more than 1 missing words or (c) more than 1 misspelt words or d) contains individual fluent word-groups with missing coherence between them.
3. Score 3: Moderate. Sentence is predominantly fluent but contains either (a) misspelt word or (b) missing word or (c) multiple occurrence of a word.
4. Score 4: Perfect. Perfectly fluent sentence without any syntactic or grammatical error.
Strictly respond in the form of JSON with the following format: {”S1”: the score, ”S2”: the score}.
Sentences: {dictionary of sentences}
On obtaining 4000 training and 700 validation samples, we finetune a Bert-base (110M parameters) to train a local discriminator.
C.4 Generator for Longer Attributes
In this section, we explore the generator’s capability to replace longer attributes.
Dataset Construction: To explore the generator’s ability on generating longer attributes, we craft a privacy dataset for ConfusionPrompt (PrivCP) with sensitive information from 7 categories. For each category, we utilize GPT-4 to generate 50 simulated user queries. Then we manually inspect and select 20 queries for necessary grammatical and semantic adjustments. Table 6 presents an example from each category and its associated risks.
In Table 8, we summarize the statistic of private attribute for strategyQA, MuSiQue, and PrivCP in terms of three metrics: i) average number of private attributes per query, ii) average length of each attribute (measured by the number of words), and iii) average attribute-to-query ratio (measured by the percentage of attribute words in a query). It can be observed that PrivCP has much longer attributes than the two standard benchmarks.
Table 8: Statistics of private attributes on three datasets.
Avg. # of attributes Avg. length of attributes Avg. attribute- to-query Ratio
strategyQA 2.18 1.74 40.88%
MuSiQue 2.82 2.72 43.20%
PrivCP 2.52 6.75 73.80%
   Avg. stands for average.
Generation Performance: The generation quality is evaluated with the following measures: i) BLEU [52] between the generated psuedo sentence and a set of referenced replacements, ii) fluency score on a scale from 1 to 4, and iii) semantic similarity scores ranging from 0 to 1. Details on the computation of fluency and similarity scores are provided in Appendix C.3.
We evaluate the generation performance on three language models in Table 9. It can be observed that while there is no significant variation of similarity score among models, Flan-T5-large has the best performance in terms of BLEU and fluency score.
Table 9: Generation performance and parameter size of various language models.
BLEU Fluency Score Similarityy Score Paratemetr Size
BART-large 0.345 2.9 0.53 406M
T5-large 0.406 3.66 0.557 737M
Flan-T5-large 0.428 3.7 0.555 783M
C.5 Accuracy of Discriminator
We assess the reliability of the local discriminator with two metric: accuracy (ACC) and mean squared error (MSE). The discriminator is trained on 4000 samples and tested on 700 samples. To validate the capability of a GPT-4 scorer, we constructe 100 samples with manual labeling and obtain the performance of GPT-4 judger on this dataset. According to Table 10, the ACC of local discriminator is above 0.7 for a 4-label classification problem, with MSE approximately 16% of the total scale (0-3). The scores given by GPT-4 judger are also close to the human labels.
Table 10: Accuracy and MSE for local discriminator and GPT-4 judger.
Local discriminator GPT-4 judger
ACC MSE ACC MSE
Metrics 0.710 0.488 0.800 0.200
   The local discriminator is tested on the validation data scored by GPT-4. GPT-4 Judger is evaluated against ground-truth data with manual label.

### v5

@@ 5.3.5 Local Model Selection
In this section, we justify the selection of decomposer and recomposer. For decomposer, we obtain the BLEU between the generated decomposition and reference provided by the dataset. For recomposer, the models are evaluated in terms of the accuracy metrics described in Section 5.3.2, using the golden decomposition and sub-answer as training and testing inputs.
Table 3 presents the experiment results on MuSiQue using models from three categories (i.e., BART, T5, and Flan-T5) of varying sizes. While Flan-T5-large exhibits the best performance for both tasks, we select BART-large as the decomposition and recomposition model due to its balanced trade-off between model size and performance.
@@
Decomposer Recomposer
@@
BART-base (139M) 0.314 0.563 0.564 0.446
BART-large (406M) 0.402 0.640 0.641 0.525
T5-base (223M) 0.343 0.581 0.580 0.484
T5-large (737M) 0.408 0.646 0.645 0.554
Flan-T5-base (248M) 0.367 0.608 0.609 0.518
Flan-T5-large (783M) 0.458 0.671 0.671 0.577
@@ 6 Discussion
ConfusionPrompt focuses on the privacy protection for textual data during privacy-preserving LLM inference. In this section, we provide further clarifications and discussions on practical issues of our framework.
6.1 Protection Capability of Privacy Model
@@
Our $(\lambda,\mu,\rho)$-privacy model establishes requirements for fabricated prompts to mitigate privacy risks. Rather than offering a worst-case privacy guarantee such as LDP, the model quantifies the criteria that a group of prompts should satisfy. Our theoretical and empirical experiments associate the privacy parameters with attack success rate. It is important to note that the privacy model should be regarded as an enhancement over the insecure status quo, rather than a substitute for formal privacy guarantees.
Alternative privacy models could be explored to enhance privacy protection in future studies. One potential direction is to analyze the distribution of sensitive attributes conditioned on the non-sensitive components, ensuring that the set of pseudo-prompts closely approximates the ground-truth distribution.
6.2 Trade-off between Privacy and Query Cost
@@
Whereas traditional perturbation-based methods trade privacy for utility, our ConfusionPrompt trades privacy for query complexity while providing accurate responses. This approach is particularly suitable for clients with high privacy demands who can accommodate these costs, such as finance and healthcare organizations. Additionally, although query costs rise linearly with the number of tokens in API services, users of chatbot services can submit unlimited requests under a fixed subscription fee, allowing for enhanced privacy protection at a manageable cost.
The query complexity could be further reduced by considering more condensed representation of the batch of pseudo-prompt (see .8). While the complexity still remains at least linearly dependent on the significance requirement $1/\mu$, future research may explore approaches to achieve a sub-linear dependence on $1/\mu$.
6.3 Generation of Pseudo-Prompts
The privacy protection of our framework relies on the quality of pseudo-prompts in effectively concealing sensitive attributes. In this work, we define the quality of fabricated prompts based on three key privacy considerations: the significance of attributes, the genuineness of the pseudo-query, and its similarity to the genuine prompt. Additionally, the generated prompts should maintain a consistent semantic structure for the non-sensitive components. The privacy requirement and assumption are satisfied by selecting filtered prompts from the generated candidates. Our approach employs a lightweight local generator to produce fabricated prompts; however, this may be insufficient, particularly when generating a large number of pseudo-prompts under stringent privacy requirements. To address this, we propose an alternative method in .9 that leverages a powerful online LLM for pseudo-prompt generation while ensuring that private attributes remain hidden from the LLM owner.
In Section 5.3.1, we present results from prompt identification attack, demonstrating the difficulty in distinguishing real prompt from pseudo-prompts under appropriate choice of privacy parameters. While more advanced discriminators may achieve higher attack success rates, developing stronger attackers is left for future work. This may involve using LLMs as discriminators and integrating domain knowledge. To further enhance privacy protection, future improvements could: (1) enhance quality assessment by incorporating additional privacy requirement, such as the diversity of pseudo-prompts, to strengthen privacy protection; (2) refine the evaluation of genuineness and similarity in experiments by considering metrics beyond fluency scores and embedding cosine similarity; (3) train the generator using adversarial training techniques and employ a set of sophisticated discriminators to select qualified pseudo-prompts, thereby improving the robustness of pseudo-prompt generation.
6.4 User Control over Sensitive Data
For privacy protection, we implement the decomposer and generator locally on the user side, granting users complete control over their sensitive data. This setup allows users to customize pseudo-prompt generation based on their privacy and ethical requirements, domain expertise, and computational resources. In particular:
  • Users can freely set the privacy parameters $(\lambda,\mu,\rho)$. For stringent privacy needs, users might: (1) assign a lower value to the significance parameter $\mu$, resulting in a larger set of pseudo-prompts; (2) set a lower similarity parameter $\lambda$ and a higher genuineness parameter $\rho$ by adjusting the filtering threshold.
  • During pseudo-prompt generation, users can filter prompts to achieve the requirement of $\lambda$ and $\rho$ through: (1) utilizing a local similarity and discrimination model (see Appendix .6), ensuring intermediate results remain on the user’s device for enhanced privacy; (2) employing another remote LLM to assess pseudo-prompt quality, offloading computational tasks to a server, which maintains privacy if the two LLM servers do not collaborate. Additionally, users can manually review and modify prompt groups before sending them to the server.
@@ 7 Conclusion
This paper proposes a private inference framework for online LLMs, termed ConfusionPrompt. We deploy three local models on the user side: (i) decomposer that maps an original prompt to a sequence of sub-prompts, (ii) generator that produces pseudo-prompts by replacing private attributes in the genuine sub-prompts, (iii) recomposer that maps the decomposed prompt-response pairs from the cloud service to the final response. Such design endows our framework with advantages over previous protocols that: (i) it can be seamlessly integrated with existing black-box LLMs where the service providers have no need to modify their already-built framework, and (ii) it achieves better privacy-utility trade-off than existing LDP-based privatization methods. We develop a $(\lambda,\mu,\rho)$-privacy model to formulate the requirement for a privacy-preserving group of prompts, and accordingly provide a complexity analysis to demonstrate the benefits of the decomposition module. Experiments show that our ConfusionPrompt outperforms local inference with open-source models, while reducing memory consumption by at least 62% compared to open-source LLMs. Furthermore, it achieves over 42% higher utility than LDP-based methods.